Sage Inc. manufactures cycling equipment. Recently, the vice president of operations of the company has requested construction of a new plant to meet the increasing demand for the company’s bikes. After a careful evaluation of the request, the board of directors has decided to raise funds for the new plant by issuing $3,454,800 of 11% term corporate bonds on March 1, 2020, due on March 1, 2035, with interest payable each March 1 and September 1, with the first interest payment on September 1st, 2020. At the time of issuance, the market interest rate for similar financial instruments is 10%.

As the controller of the company, determine the selling price of the bonds. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Selling price of the bonds
Answer #1

Face Value of bond is missing in the problem, so i am assuming face vakue of bond to be $1000.

Slae price of bonds will be sum of present value of coupons adding with present value of face value of bond.

So present value of future coupon payable twice a year will be calculate with interest rate of 6.5% (instead of 11% rate will be 6.5% as coupon payable monthly) for a time period of 30 times (instead of 15 years time will be 30 years because coupon is payable monthly).

Present value of coupon= 65*present value factor with rate of 5% for 30 years

=65*15.372=$999.18

Present Vakue of Face Vakue=1000*0.2314=$231.4

Total Sale Price of Bond=$999.18+$231.4=$1231.58

Final Sale Price=$1231.

Present value of annuity of $1 for 30 years with Rate of 5%= 15.372

Oresent value of $1 with rate of 5% for 30 years will be =0.2314